A motorist traveling with a constant speed of 20.0 m/s passes a school-crossing corner, where the speed limit is 15.0 m/s. Just as the motorist passes, a police officer on a motorcycle at the corner starts off in pursuit with constant acceleration of 5.0 m/s^2. What is the total distance traveled by the motorist

1
Expert's answer
2022-03-03T12:04:47-0500
"d_M=v_0t=20t"

"d_C=at^2\/2=2.5t^2"

"d_M=d_C"

"20t=2.5t^2"

"t=8\:\rm s"

The total distance traveled by the motorist

"d_M=20*8=160\:\rm m"